OpenStudy (anonymous):

Okay so I have to plot points, which is rather easy to do... So I go to mathway.com and put in -2x-3y=-7 , right? Then I click the option to find three ordered pairs. But then I got things like (0,7/3) , (1,5/3), and (2,1) Well the last one was easy to plot. But the first two .... Errr. I'm on a program doing a test and it'll only let me plot whole numbers like (2,1) If someone could help me or help explain how to plot those, or something that'd be great ! xoxo

OpenStudy (anonymous):

Geogebra is better!

OpenStudy (phi):

if you put the equation into y = mx + b form you get \[ y= -\frac{2}{3} x + \frac{7}{3} \] as you noticed (2,1) works: y = -(2/3)*2 + 7/3 = -4/3 +7/3 = 3/3 = 1 next, notice the slope is -2/3 change in y is -2 , change in x is 3 that means if you add 3 to the x , the y goes down by 2 (i.e. add -2)

OpenStudy (phi):

once you find a nice point such as (2,1) you can add 3 to the x and -2 to the y to get another nice point

OpenStudy (anonymous):

Use Geogebra because it just worked for me and is easier.